# A Calculus of Inheritance

### v4

@@ Appendix B Multi-Path Self-Reference in Other Systems
Appendix C Well-Definedness of the Semantic Functions
This appendix proves that the mutually recursive equations (1)–(6) have a unique least fixed point, and therefore the semantics of Section 3 is well-defined.
C.1. The Product Lattice
Fix an AST with its primitive functions $\mathtt{defines}$ and $\mathtt{inherits}$. Let $\mathcal{L}$ denote the set of all labels and $\mathcal{P}$ the set of all paths (finite sequences of labels). Define the following domains, one for each semantic function:
$D_{\mathtt{properties}}$ $=\mathcal{P}\to\mathcal{P}(\mathcal{L})$ (sets of labels)
$D_{\mathtt{supers}}$ $=\mathcal{P}\to\mathcal{P}(\mathcal{P}\times\mathcal{P})$ (sets of site–override pairs)
$D_{\mathtt{overrides}}$ $=\mathcal{P}\to\mathcal{P}(\mathcal{P})$ (sets of paths)
$D_{\mathtt{bases}}$ $=\mathcal{P}\to\mathcal{P}(\mathcal{P})$ (sets of paths)
$D_{\mathtt{resolve}}$ $=(\mathcal{P}\times\mathcal{P}\times\mathbb{N}\times\mathcal{L}^{*})\to\mathcal{P}(\mathcal{P})$ (sets of paths)
$D_{\mathtt{this}}$ $=(\mathcal{P}(\mathcal{P})\times\mathcal{P}\times\mathbb{N})\to\mathcal{P}(\mathcal{P})$ (sets of paths)
Each $D_{i}$ is a complete lattice under the pointwise subset ordering: $f\sqsubseteq g$ iff $f(x)\subseteq g(x)$ for all $x$. The product $D=D_{\mathtt{properties}}\times D_{\mathtt{supers}}\times D_{\mathtt{overrides}}\times D_{\mathtt{bases}}\times D_{\mathtt{resolve}}\times D_{\mathtt{this}}$ is a complete lattice under the componentwise ordering, with bottom element $\bot=(\lambda x.\,\varnothing,\ldots,\lambda x.\,\varnothing)$.
C.2. The Operator
Equations (1)–(6) define a function $F:D\to D$ that takes a tuple of candidate interpretations $(\mathtt{properties},\mathtt{supers},\mathtt{overrides},\mathtt{bases},\mathtt{resolve},\mathtt{this})$ and produces a new tuple by evaluating the right-hand side of each equation. The transitive closure $\mathtt{bases}^{*}$ appearing in equation (2) is derived from $\mathtt{bases}$: for any fixed $\mathtt{bases}$, the set $\mathtt{bases}^{*}(p)$ is defined as the least fixed point of the monotone operator $X\mapsto\{p\}\cup\bigcup_{q\in X}\mathtt{bases}(q)$ on $\mathcal{P}(\mathcal{P})$. If $\mathtt{bases}\sqsubseteq\mathtt{bases}^{\prime}$ (pointwise), then $\mathtt{bases}^{*}\sqsubseteq\mathtt{bases}^{\prime*}$ (pointwise), since enlarging the one-step relation can only enlarge the transitive closure.
The $\mathtt{this}$ function involves a recursion on the de Bruijn index $n$. Since $n$ decreases by one at each step and is a non-negative integer, this recursion terminates for any fixed interpretation of $\mathtt{supers}$. The result of $\mathtt{this}(S,p_{\mathrm{def}},n)$ for all valid $(S,p_{\mathrm{def}},n)$ is therefore well-defined as a function of $\mathtt{supers}$, and it remains to show that it is monotone in $\mathtt{supers}$.
C.3. Monotonicity
Proposition C.1.
The operator $F$ is monotone: if $\vec{d}\sqsubseteq\vec{d}^{\prime}$ in $D$, then $F(\vec{d})\sqsubseteq F(\vec{d}^{\prime})$.
Proof.
It suffices to show that each component of $F$ is monotone in every argument, i.e., that enlarging any of the input interpretations (in the $\subseteq$ ordering) can only enlarge the output. We verify this by inspecting the right-hand side of each equation and confirming that every occurrence of a mutually defined function appears in a positive position—as the domain of an existential quantifier in a set comprehension (Aczel, 1977).
Equation (1): $\mathtt{properties}$
$\mathtt{properties}(p)=\bigl\{\,\ell\bigm|(\_,p_{\mathrm{override}})\in\mathtt{supers}(p),\;\ell\in\mathtt{defines}(p_{\mathrm{override}})\,\bigr\}.$
The function $\mathtt{supers}$ appears as the set being iterated over ($\in\mathtt{supers}(p)$). The predicate $\ell\in\mathtt{defines}(\cdot)$ involves only the AST primitive $\mathtt{defines}$, not any mutually defined function. Enlarging $\mathtt{supers}(p)$ can only add pairs, hence can only add labels to the result. Positive in $\mathtt{supers}$.
Equation (2): $\mathtt{supers}$
$\mathtt{supers}(p)=\bigl\{\,(\mathtt{init}(p_{\mathrm{base}}),p_{\mathrm{override}})\bigm|p_{\mathrm{base}}\in\mathtt{bases}^{*}(p),\;p_{\mathrm{override}}\in\mathtt{overrides}(p_{\mathrm{base}})\,\bigr\}$. Both $\mathtt{bases}^{*}$ and $\mathtt{overrides}$ appear as sets being iterated over. As noted above, $\mathtt{bases}^{*}$ is monotone in $\mathtt{bases}$. Enlarging either $\mathtt{bases}$ or $\mathtt{overrides}$ can only add elements to the result. Positive in $\mathtt{bases}$ and $\mathtt{overrides}$.
Equation (3): $\mathtt{overrides}$
For $p\neq()$:
$\mathtt{overrides}(p)=\{p\}\cup\bigl\{\,p_{\mathrm{branch}}\mathbin{\triangleright}\mathtt{last}(p)\bigm|{}\\ (\_,p_{\mathrm{branch}})\in\mathtt{supers}(\mathtt{init}(p)),\;\mathtt{last}(p)\in\mathtt{defines}(p_{\mathrm{branch}})\,\bigr\}.$
The function $\mathtt{supers}$ appears as the set being iterated over. The filtering condition $\mathtt{last}(p)\in\mathtt{defines}(\cdot)$ involves only the AST primitive. Enlarging $\mathtt{supers}$ can only add branches, hence can only add paths to the result. Positive in $\mathtt{supers}$.
Equation (4): $\mathtt{bases}$
$\mathtt{bases}(p)=\bigl\{\,p_{\mathrm{target}}\bigm|p_{\mathrm{override}}\in\mathtt{overrides}(p),\\ (n,\ell_{\mathrm{down},*})\in\mathtt{inherits}(p_{\mathrm{override}}),\;p_{\mathrm{target}}\in\mathtt{resolve}(\ldots)\,\bigr\}.$
Both $\mathtt{overrides}$ and $\mathtt{resolve}$ appear as sets being iterated over. The function $\mathtt{inherits}$ is an AST primitive. Enlarging either $\mathtt{overrides}$ or $\mathtt{resolve}$ can only add targets to the result. Positive in $\mathtt{overrides}$ and $\mathtt{resolve}$.
Equation (5): $\mathtt{resolve}$
$\mathtt{resolve}(\ldots)=\bigl\{\,p_{\mathrm{current}}\mathbin{\triangleright}\ell_{\mathrm{down},1}\mathbin{\triangleright}\cdots\bigm|p_{\mathrm{current}}\in\mathtt{this}(\{p_{\mathrm{site}}\},\mathtt{init}(p_{\mathrm{def}}),n)\,\bigr\}.$
The function $\mathtt{this}$ appears as the set being iterated over. Enlarging $\mathtt{this}$ can only add paths to the result. Positive in $\mathtt{this}$.
Equation (6): $\mathtt{this}$
For $n>0$: $\mathtt{this}(S,p_{\mathrm{def}},n)=\mathtt{this}(S^{\prime},\mathtt{init}(p_{\mathrm{def}}),n-1)$ where $S^{\prime}=\bigl\{\,p_{\mathrm{site}}\bigm|p_{\mathrm{current}}\in S,\;(p_{\mathrm{site}},p_{\mathrm{override}})\in\mathtt{supers}(p_{\mathrm{current}}),\;p_{\mathrm{override}}=p_{\mathrm{def}}\,\bigr\}$. The function $\mathtt{supers}$ appears as the set being iterated over. The filtering condition $p_{\mathrm{override}}=p_{\mathrm{def}}$ is an equality test on path data, not involving any mutually defined function. Enlarging $\mathtt{supers}$ can only enlarge $S^{\prime}$, and by induction on $n$ (the base case $n=0$ returns $S$ unchanged), the recursive call is monotone in $S^{\prime}$. Positive in $\mathtt{supers}$.
In summary, every mutually defined function appears on the right-hand side of every equation only inside the pattern $x\in f(\ldots)$ within a set comprehension, i.e., as the domain of an existential quantifier. No complement, set difference, or negated membership test appears in any equation. Therefore $F$ is a monotone operator on the complete lattice $D$. ∎
C.4. Existence of the Least Fixed Point
Theorem C.2.
The semantic functions $\mathtt{properties}$, $\mathtt{supers}$, $\mathtt{overrides}$, $\mathtt{bases}$, $\mathtt{resolve}$, $\mathtt{this}$ are well-defined as the least fixed point of $F$.
Proof.
By Proposition C.1, $F$ is monotone on the complete lattice $D$. By the Knaster–Tarski theorem (Tarski, 1955), $F$ has a least fixed point $\vec{d}^{*}=\bigsqcup_{n\geq 0}F^{n}(\bot)$, where $\bot$ assigns the empty set to every input. The components of $\vec{d}^{*}$ are the semantic functions of Section 3.
Concretely, the least fixed point is the union of the finite approximation chain: $\bot\sqsubseteq F(\bot)\sqsubseteq F^{2}(\bot)\sqsubseteq\cdots$. A membership $\ell\in\mathtt{properties}(p)$ holds in $\vec{d}^{*}$ if and only if it appears at some finite stage $F^{k}(\bot)$, which corresponds to a finite derivation tree using the equations. When no finite derivation exists, the membership is absent from the least fixed point and the query is undefined, the analogue of divergence. ∎
Appendix D Böhm Tree Correspondence: Proofs
This appendix contains the full proofs of Adequacy (Theorem 4.2) and Full Abstraction (Theorem 4.3) stated in Section 4.1.
D.1. Böhm Trees
Böhm tree equivalence $\mathrm{BT}(M)=\mathrm{BT}(N)$ is the standard observational equivalence of the lazy $\lambda$-calculus (Barendregt, 1984; Abramsky and Ong, 1993).
D.2. Path Encoding
  • A tail call $V_{1}\;V_{2}$ translates to a record with own properties $\{\_\mathrm{tail},\mathrm{result}\}$, where $\mathrm{tail}$ holds the encapsulated application and $\mathrm{result}$ projects $\mathrm{tail}.\mathrm{result}$.
D.3. Single-Path Lemma
Lemma D.1 (Single Path).
Proof.
By induction on the structure of the ANF term.
Case $\mathcal{T}(V_{1}\;V_{2})=\{\_\mathrm{tail}\mapsto\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\},\;\mathrm{result}\mapsto\_\mathrm{tail}.\mathrm{result}\}$
Identical to the $\mathbf{let}$-binding case: the outer record has own properties $\{\_\mathrm{tail},\mathrm{result}\}$ with no inheritance sources at the outer level, and the $\mathrm{tail}$ subtree contains exactly one inheritance source.
Case $\mathcal{T}(x)=\uparrow^{\!n}.\mathrm{argument}$
This is a reference, not a record. The $\mathtt{this}$ function walks up $n$ steps from the enclosing scope. By the inductive hypothesis on the enclosing translations, each step of $\mathtt{this}$ encounters exactly one matching pair in $\mathtt{supers}$.
The key structural invariant is that the translation $\mathcal{T}$ never produces two independent inheritance sources at the same scope level that both define the same label. In every case, application inheritances are encapsulated inside a named property ($x$ or $\mathrm{tail}$), and each such encapsulated inheritance has exactly one inheritance source. This prevents diamond inheritance and ensures single-path resolution throughout. ∎
D.4. Substitution Lemma
Lemma D.2 (Substitution).
where $M[V/x]$ is the usual capture-avoiding substitution.
Proof.
By structural induction on $M$.
Base case: $M=x$
Then $\mathcal{T}(x)=\uparrow^{\!n}.\mathrm{argument}$ (where $n$ is the de Bruijn index of $x$), and $C\mathbin{\triangleright}\mathrm{result}$ contains this reference. In the inherited tree $C$, $\mathtt{resolve}$ and $\mathtt{this}$ resolve the reference by navigating from the reference’s enclosing scope up $n$ steps to the binding $\lambda$, where the inheritance $\{\mathcal{T}(\lambda x.\,x),\;\mathrm{argument}\mapsto\mathcal{T}(V)\}$ inherits into the $\mathrm{argument}$ slot, providing $\mathcal{T}(V)$. Since $C$ is the $\mathrm{tail}$ subtree of $\mathcal{T}\bigl((\lambda x.\,M)\;V\bigr)$, which is a closed ANF term, Lemma D.1 applies and $\mathtt{this}$ finds exactly one path, so $\mathtt{resolve}$ returns $\mathcal{T}(V)$’s subtree. Since $M[V/x]=V$, we have $\mathcal{T}(M[V/x])=\mathcal{T}(V)$, and the properties coincide.
Base case: $M=y$ where $y\neq x$
[…]

### v6

@@ Appendix D Multi-Path Self-Reference in Other Systems
Appendix E First-Order Semantics of the ANF Lazy $\lambda$-Calculus
This appendix presents a first-order semantics of the lazy $\lambda$-calculus in A-normal form (ANF). The semantics is given by mutually recursive set-comprehension equations whose well-definedness follows from the same termination argument as in Appendix A. The central idea is that function application is modeled as tree grafting: applying a function record to an argument splices the function’s subtree into the call site, and variable references navigate the resulting tree structure.
ANF is essential. Traditional $\lambda$-calculus syntax allows anonymous subexpressions such as (fx)(f\;x) in (fx)y(f\;x)\;y that have no label and cannot form paths. ANF restores path-addressability by binding every intermediate result to a name, so that every node in the evaluation tree has a unique path address.
Syntax
Let $M$ denote a computation and $V$ a value.
[…]
The grammar includes both application let-bindings and value let-bindings (the latter bind a name to a value without application). Every $\lambda$-term can be mechanically converted to ANF by naming intermediate results; ANF is neither a restriction nor an extension of the $\lambda$-calculus. Variables use de Bruijn indices: $\uparrow^{\!n}$ counts the number of enclosing scope levels (record braces) between the reference and its binder, following the same convention as inheritance-calculus. In particular, a $\lambda$-bound variable that crosses $k$ intervening $\mathbf{let}$-binding scopes and $n$ intervening $\lambda$-abstraction scopes is written $\uparrow^{\!n+k}$, not $\uparrow^{\!n}$. Let-bound variables are referenced by their binding name $x$.
Evaluation tree
An evaluation tree has a single abstraction: every node can serve as function, argument, binding, and result at the same time. A $\mathtt{Path}$ is a sequence of labels $(\ell_{1},\ldots,\ell_{k})$ identifying a position in the mixin tree. We write $p$ for a path, $()$ for the root, $p\mathbin{\triangleright}\ell$ for extension, $\mathtt{init}(p)$ for the parent path, and $\mathtt{last}(p)$ for the final label.
AST
An ANF term is parsed into an AST that provides two functions at each path $p$:
  • $\mathtt{children}(p)$: the set of labels locally defined at $p$.
  • $\mathtt{refs}(p)$: the set of reference pairs $(n,\;\ell_{*})$, where $n$ is a de Bruijn index and $\ell_{*}$ is a list of downward projections.
Each ANF construct populates these as follows.
An abstraction $\lambda x.\,M$ at path $p$:
$\mathtt{children}(p)$ $=\{\mathrm{argument},\;\mathrm{result}\},$ $\mathtt{refs}(p)$ $=\varnothing.$
The subtree at $p\mathbin{\triangleright}\mathrm{result}$ is determined by $M$; the subtree at $p\mathbin{\triangleright}\mathrm{argument}$ is empty ($\mathtt{children}=\varnothing$, $\mathtt{refs}=\varnothing$).
A let-binding $\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M$ at path $p$:
$\mathtt{children}(p)$ $=\{x,\;\mathrm{result}\},$ $\mathtt{refs}(p)$ $=\varnothing.$
The subtree at $p\mathbin{\triangleright}\mathrm{result}$ is determined by $M$. The subtree at $p\mathbin{\triangleright}x$ represents the application $V_{1}\;V_{2}$ and is determined by the form of $V_{1}$:
  • If $V_{1}$ is a variable (either $\uparrow^{\!m}$ or a let-bound name), then
    $\mathtt{children}(p\mathbin{\triangleright}x)$ $=\{\mathrm{argument}\},$ $\mathtt{refs}(p\mathbin{\triangleright}x)$ $=\{(n_{1},\;\ell_{*})\},$
    where the reference $(n_{1},\;\ell_{*})$ encodes the translation of $V_{1}$ (see below).
  • If $V_{1}=\lambda y.\,M^{\prime}$, the $\lambda$-literal is inlined:
    $\mathtt{children}(p\mathbin{\triangleright}x)$ $=\{\mathrm{argument},\;\mathrm{result}\},$ $\mathtt{refs}(p\mathbin{\triangleright}x)$ $=\varnothing,$
    and the subtree at $p\mathbin{\triangleright}x\mathbin{\triangleright}\mathrm{result}$ is determined by $M^{\prime}$.
In both cases the subtree at $p\mathbin{\triangleright}x\mathbin{\triangleright}\mathrm{argument}$ is determined by $V_{2}$.
A value let-binding $\mathbf{let}\;x=V\;\mathbf{in}\;M$ at path $p$:
$\mathtt{children}(p)$ $=\{x,\;\mathrm{result}\},$ $\mathtt{refs}(p)$ $=\varnothing.$
The subtree at $p\mathbin{\triangleright}\mathrm{result}$ is determined by $M$. The subtree at $p\mathbin{\triangleright}x$ wraps the value in a $\mathrm{result}$ scope:
$\mathtt{children}(p\mathbin{\triangleright}x)$ $=\{\mathrm{result}\},$ $\mathtt{refs}(p\mathbin{\triangleright}x)$ $=\varnothing.$
The subtree at $p\mathbin{\triangleright}x\mathbin{\triangleright}\mathrm{result}$ is determined by $V$. The accessor $x.\mathrm{result}$ thus yields the translated value, matching the convention for application let-bindings.
A tail call $V_{1}\;V_{2}$ at path $p$:
$\mathtt{children}(p)$ $=\{\mathrm{tailCall},\;\mathrm{result}\},$ $\mathtt{refs}(p)$ $=\varnothing.$
The subtree at $p\mathbin{\triangleright}\mathrm{tailCall}$ is analogous to the application node $p\mathbin{\triangleright}x$ above: if $V_{1}$ is a variable, $\mathtt{children}=\{\mathrm{argument}\}$ with $\mathtt{refs}=\{(n_{1},\;\ell_{*})\}$; if $V_{1}$ is a $\lambda$-literal, $\mathtt{children}=\{\mathrm{argument},\;\mathrm{result}\}$ with $\mathtt{refs}=\varnothing$ and the $\lambda$-body inlined. The subtree at $p\mathbin{\triangleright}\mathrm{result}$ is a forwarding reference:
$\mathtt{children}(p\mathbin{\triangleright}\mathrm{result})$ $=\varnothing,$ $\mathtt{refs}(p\mathbin{\triangleright}\mathrm{result})$ $=\{(0,\;(\mathrm{tailCall},\;\mathrm{result}))\}.$
Here $(0,\;(\mathrm{tailCall},\;\mathrm{result}))$ navigates to the enclosing scope ($p$ itself, since $n=0$ from $p\mathbin{\triangleright}\mathrm{result}$’s perspective), then projects $\mathrm{tailCall}.\mathrm{result}$.
A $\lambda$-bound variable $\uparrow^{\!n}$ used as an inheritance source at an application node $p$ contributes $(n,\;(\mathrm{argument}))$ to $\mathtt{refs}(p)$: navigate up $n$ scope levels, then project $\mathrm{argument}$. The same applies when $\uparrow^{\!n}$ occurs as a bare value in tail position or as an argument value, contributing the reference to $\mathtt{refs}$ at the corresponding path.
A let-bound variable $x$ at path $p$ is resolved lexically: find the nearest ancestor $p^{\prime}$ of $p$ such that $x\in\mathtt{children}(p^{\prime})$, let $n=|p|-|p^{\prime}|-1$, and contribute $(n,\;(x,\;\mathrm{result}))$ to $\mathtt{refs}(p)$, projecting $x.\mathrm{result}$ from the enclosing scope.
Note that since the de Bruijn index $n$ counts all scope levels (not just $\lambda$-abstractions), it accounts for intervening $\mathbf{let}$-binding and tail-call scopes. For example, in $\lambda.\,\lambda.\,\mathbf{let}\;r=\uparrow^{\!2}\;\uparrow^{\!0}\;\mathbf{in}\;r$, the outer-$\lambda$ variable has de Bruijn index $2$ (crossing the $\mathbf{let}$-scope and the inner-$\lambda$ scope), not $1$.
Both $\mathtt{children}$ and $\mathtt{refs}$ are pure data extracted from the AST; they are not functions of runtime state.
Labels
The labels observable at a path are those locally defined at any graft of any transitive callee:
(7) $\mathtt{labels}(p)=\bigl\{\;\ell\;\big|\;p_{\mathrm{step}}\in\mathtt{callee}^{*}(p),\;p_{\mathrm{graft}}\in\mathtt{grafts}(p_{\mathrm{step}}),\;\ell\in\mathtt{children}(p_{\mathrm{graft}})\;\bigr\}$
Grafts
When a function is applied, its subtree is grafted at the call site. A child label $\ell$ of the callee also becomes a child of the caller; the child at the callee’s $\ell$ position and the child at the caller’s $\ell$ position (if any) share the same identity. The grafts of $p$ collect all such identity-sharing positions:
(8) $\mathtt{grafts}(p)=\begin{cases}\{p\}&\text{if }p=()\\[6.0pt] \{p\}\;\cup\;\left\{\;p_{\mathrm{callee}}\mathbin{\triangleright}\mathtt{last}(p)\;\left|\;\begin{aligned} &p_{\mathrm{callee}}\in\mathtt{callee}^{*}(\mathtt{init}(p)),\\ &\text{s.t.}\;\mathtt{last}(p)\in\mathtt{children}(p_{\mathrm{callee}})\end{aligned}\right.\right\}&\text{if }p\neq()\end{cases}$
Here $\mathtt{callee}^{*}$ denotes the reflexive-transitive closure of $\mathtt{callee}$, defined next.
Callee
The callee of a path $p$ is the set of paths that $p$ is grafted from. Concretely, $\mathtt{callee}$ resolves every reference in $p$’s grafts one step:
[…]
where $\ell_{*}=(\ell_{1},\ldots,\ell_{k})$. The $\mathtt{init}(p_{\mathrm{graft}})$ passes the reference’s enclosing scope to $\mathtt{scope}$, matching the convention of reference resolution (5), which calls $\mathtt{this}(\{p_{\mathrm{site}}\},\;\mathtt{init}(p_{\mathrm{def}}),\;n)$.
Scope resolution
$\mathtt{scope}$ navigates upward through the evaluation tree to find the binding site of a de Bruijn reference. Each ANF construct ($\lambda$-abstraction, $\mathbf{let}$-binding, or tail call) introduces one scope level; crossing a scope boundary means finding the call site that grafted the current subtree.
$\mathtt{scope}$ takes three arguments: an evaluation-site path $p_{\mathrm{site}}$ (where the grafted subtree lives), a definition-site path $p_{\mathrm{def}}$ (the position in the original AST), and a de Bruijn index $n$.
(10) $\mathtt{scope}(p_{\mathrm{site}},\;p_{\mathrm{def}},\;n)=\begin{cases}p_{\mathrm{site}}&\text{if }n=0\\[6.0pt] \mathtt{scope}\!\left(p_{\mathrm{caller}},\;\mathtt{init}(p_{\mathrm{def}}),\;n-1\right)&\text{if }n>0\end{cases}$
where $p_{\mathrm{caller}}$ is the unique path satisfying $(p_{\mathrm{caller}},\;p_{\mathrm{def}})\in\mathtt{callee}^{\mathrm{ctx}}(p_{\mathrm{site}})$, and $\mathtt{callee}^{\mathrm{ctx}}$ pairs each element of $\mathtt{callee}^{*}(p_{\mathrm{site}})$ with its provenance:
(11) $\mathtt{callee}^{\mathrm{ctx}}(p)=\bigl\{\;(\mathtt{init}(p_{\mathrm{step}}),\;p_{\mathrm{graft}})\;\big|\;p_{\mathrm{step}}\in\mathtt{callee}^{*}(p),\;p_{\mathrm{graft}}\in\mathtt{grafts}(p_{\mathrm{step}})\;\bigr\}$
At each step, $\mathtt{init}$ shortens $p_{\mathrm{def}}$ by one label and $n$ decreases by one. Since $n$ is a nonnegative integer, the recursion terminates.
This completes the chain of definitions needed to compute $\mathtt{labels}(p)$.
Observational fixpoint semantics
The definitions above constitute the complete semantics. Computation arises from querying progressively deeper paths in a lazily constructed tree: the observer drives computation by choosing which path to inspect, and the mutually recursive definitions unfold on demand.
Formally, the mutually recursive equations (7)–(11) define set membership recursively: $\ell\in\mathtt{labels}(p)$ holds if and only if it can be established by a finite chain of applications of these equations. Well-definedness follows from the same termination argument as for the mixin tree equations (Appendix A): if the recursive evaluation terminates, the reachable dependency subgraph is finite and acyclic, and the equations have a unique solution. When no finite derivation exists for a given query, the recursive evaluation does not terminate, analogous to divergence.
Convergence
An ANF term $M$ converges, written $M{\Downarrow}$, if there exists $n\geq 0$ such that
$\{\mathrm{argument},\,\mathrm{result}\}\;\subseteq\;\mathtt{labels}\!\bigl(\,\underbrace{\mathrm{result}\mathbin{\triangleright}\cdots\mathbin{\triangleright}\mathrm{result}}_{n}\,\bigr)$
in the semantics defined by the equations above. This corresponds exactly to the Böhm tree (Barendregt, 1984) of the original term: $M$ converges iff it has a head normal form (Theorem 4.2), and two terms are equivalent under this semantics iff they have the same Böhm tree (Theorem 4.3). The proofs, given in Appendix B, apply verbatim: each equation above is the specialization of the corresponding mixin tree equation (Section 3) to terms in the image of the translation $\mathcal{T}$ (Section 4), where the Single-Path Lemma (Lemma B.1) ensures that the set-valued $\mathtt{this}$ function (equation 6) degenerates to the single-valued $\mathtt{scope}$ (equation 10).
Appendix F Encoding Datalog in Inheritance-Calculus
Section 6.1 observed that inheritance-calculus natively produces the relational semantics of logic programming. This appendix makes that observation precise by giving a systematic encoding of Datalog into inheritance-calculus. The encoding uses no extensions to the calculus; it is expressible in the three primitives of Section 2. All examples in this appendix have executable MIXINv2 implementations (Yang, 2026). There are three ingredients to keep separate. First, relations are stored as tries, so tuple lookup is indexed by prefixes rather than by scanning a flat set of facts. Second, the control skeleton is written in the visitor-pattern style already used in Section 5; from a functional-programming viewpoint, this is the same role played by Scott-encoded case analysis. Third, inheritance union and tabled evaluation give the resulting program its relational, least-fixed-point semantics. The appendix proceeds in that order: data representation first, then control flow, then recursion.
F.1. Relations as Tries
A binary relation over a finite domain $D=\{a,b,c,\ldots\}$ is represented as a trie: a mixin tree whose paths encode tuples. The trie is built from three base scopes: $\mathrm{Relation}$ (abstract base), $\mathrm{Cons}$ (non-empty node), and $\mathrm{Nil}$ (terminator). Each constant $d\in D$ inherits from $\mathrm{Cons}$ and carries a $\mathrm{TailMap}$ that provides per-constant sub-tries:
$\mathrm{Relation}\mapsto\{\},\quad\mathrm{Cons}\mapsto\{\mathrm{Relation},\;\mathrm{Tail}\mapsto\{\mathrm{Relation}\}\},\quad\mathrm{Nil}\mapsto\{\mathrm{Relation}\}$
$d\mapsto\{\mathrm{Cons},\;\mathrm{Tail}\mapsto\{\},\;\mathrm{TailMap}\mapsto\{d\mapsto\{\mathrm{Tail}\}\}\}$
A fact $R(d_{1},d_{2})$ is encoded as a trie path of depth 2, with a $\mathrm{Nil}$ terminator. The per-constant sub-tries are accessed via $\mathrm{TailMap}.d$:
$R(a,b)\;\leadsto\quad R\mapsto\{a,\;\mathrm{TailMap}\mapsto\{a\mapsto\{b,\;\mathrm{TailMap}\mapsto\{b\mapsto\{\mathrm{Nil}\}\}\}\}\}$
Multiple facts for the same relation are accumulated by inheritance (trie union). Adding $R(b,c)$:
$R\mapsto\{a,\;b,\;\mathrm{TailMap}\mapsto\{$ $a\mapsto\{b,\;\mathrm{TailMap}\mapsto\{b\mapsto\{\mathrm{Nil}\}\}\},$
$b\mapsto\{c,\;\mathrm{TailMap}\mapsto\{c\mapsto\{\mathrm{Nil}\}\}\}\}\}$
The point of the trie representation is not merely that it is a tree-shaped encoding. It also provides the indexing structure used by the subsequent joins: once a constant $d$ has been selected, the encoding moves directly to $\mathrm{TailMap}.d$ and continues from the matching sub-trie. This is why the appendix speaks about tries rather than about an arbitrary encoding of finite sets of tuples.
F.2. Infrastructure
Each constant $d$ carries four groups of boilerplate definitions used by the encoding. These play a role analogous to derived type-class instances in Haskell: each constant must provide a fixed set of properties that the encoding’s generic patterns rely on. Inheritance-calculus has no built-in derivation mechanism, so the boilerplate is written out manually; however, the amount of boilerplate per constant is $O(1)$ and does not grow with the domain size $|D|$, so adding a new constant does not require modifying any existing scope, preserving the expression-problem property.
Acceptance:
The visitor-pattern infrastructure from Section 5. Operationally, this is the object-oriented spelling of a case split on the current constant. Each constant $d$ defines a $\mathrm{VisitorMap}$ containing a branch keyed by $d$, a $\mathrm{Visitor}$ with a $\mathrm{Visited}$ slot for per-constant results, and an $\mathrm{Accepted}$ property that delegates to the matching branch’s $\mathrm{Visited}$ value. This is the dispatch mechanism: given a scope that may contain multiple constants, $\mathrm{Acceptance}$ selects the branch corresponding to each constant present and collects the per-constant $\mathrm{Visited}$ results.
WithVisitorTail:
Maps each constant $d$ to its sub-trie entry $\mathrm{TailMap}.d$, labeled as $\mathrm{VisitorTail}$. This provides the right-hand (visitor) side of a join: when two relations’ $\mathrm{VisitorMap}$ entries are composed, $\mathrm{VisitorTail}$ carries the joined relation’s row data for the matched constant.
WithVisiteeTail:
Maps each constant $d$ to its sub-trie entry $\mathrm{TailMap}.d$, labeled as $\mathrm{VisiteeTail}$. This provides the left-hand (visitee) side: the relation being iterated. Together, $\mathrm{WithVisitorTail}$ and $\mathrm{WithVisiteeTail}$ enable value-level joins where only constants present in both relations produce output.
Replacement:
Given a $\mathrm{NewTail}$ value, produces a $\mathrm{Replaced}$ scope that inherits the label $d$ but substitutes $\mathrm{TailMap}.d$ with $\mathrm{NewTail}$. This is the output construction mechanism: it builds output tuples by substituting new sub-tries into constant wrappers. Crucially, $\mathrm{Replacement}$ is defined on the per-constant $\mathrm{TailMap}$ entry (not on the merged scope), ensuring that output construction operates on individual constants rather than their union.
As a running example, we use the executable two-hop encoding $\mathrm{twoHop}(X,Z)\mathrel{{:}\!{-}}\mathrm{edge}(X,Y),\;\mathrm{edge}(Y,Z)$ from RelationalTwoHopRepeatedVisitor.mixin.yaml (Yang, 2026). Its top-level scopes are named after the stages of the encoding: _XAcceptance iterates the first column, _YAcceptance performs the shared-variable join, _NilAcceptance0 confirms the end of the first fact, _ZAcceptance iterates the output column, and _NilAcceptance1 triggers the final Replacement chain. The schematic encodings below should be read as abstractions of that executable MIXINv2 file. If one temporarily ignores relational accumulation and reads each dispatch as an ordinary case split, the control flow is: choose $X$ from the first relation, choose a matching $Y$, confirm that the first tuple is complete, choose $Z$ from the second relation’s matching row, confirm that the second tuple is complete, and emit $(X,Z)$. The rest of the appendix explains how that familiar single-valued control skeleton is realised by visitor dispatch over tries and then lifted to relational semantics by inheritance union.
F.3. Encoding Rules
A Datalog rule $H(\bar{X})\mathrel{{:}\!{-}}B_{1}(\bar{Y}_{1}),\ldots,B_{n}(\bar{Y}_{n})$ is encoded by nesting three mechanisms inside that control skeleton:
  1. (1)
    Column iteration. For each column position, an $\mathrm{Acceptance}$ dispatch iterates over the constants present. The $\mathrm{VisitorMap}$ is populated from $\mathrm{WithVisiteeTail}$, so each matched constant’s $\mathrm{Visitor}$ receives $\mathrm{VisiteeTail}$: the per-constant sub-trie at that position. The result of processing each constant is placed in $\mathrm{Visited}$, and $\mathrm{Accepted}$ collects all per-constant results. (Details in Sections F.4, F.5, and F.6.)
  2. (2)
    Join. For each shared variable between body atoms, the inner $\mathrm{Acceptance}$’s $\mathrm{VisitorMap}$ inherits from both $\mathrm{WithVisitorTail}$ (from the joined relation) and $\mathrm{WithVisiteeTail}$ (from the iterated relation). A constant $d$ produces a $\mathrm{Visitor}$ entry only if $d$ appears in both relations, achieving a value-level join. The $\mathrm{Visitor}$ then receives both $\mathrm{VisiteeTail}$ and $\mathrm{VisitorTail}$, providing access to both relations’ data for the matched constant.
  3. (3)
    Output construction. Inside the innermost $\mathrm{Nil}$ confirmation (verifying that each body atom’s tuple is complete), chain $\mathrm{Replacement}$ operations from the innermost output column outward, each setting $\mathrm{NewTail}$ to the previous $\mathrm{Replaced}$ value, terminated by $\mathrm{Nil}$. The final $\mathrm{Replaced}$ value is assigned to $\mathrm{Visited}$.
F.4. Variable Join
Consider the simplest join: $\mathrm{composed}(X,Z)\mathrel{{:}\!{-}}\alpha(X,Y),\;\beta(Y,Z)$. The executable self-join instance of this pattern is RelationalTwoHopRepeatedVisitor.mixin.yaml (Yang, 2026), whose scope names are shown in parentheses below. This is the right place to see the division of labour from the section introduction: visitor dispatch provides the control flow, the trie provides direct access to the matching tails, and $\mathtt{supers}$ accumulates the outputs produced by each successful match.
[…]